OpenStudy (anonymous):

e^e^x=10 solve for x

myininaya (myininaya):

\[e^{e^{x}}=10\]

myininaya (myininaya):

Take ln( ) of both sides

myininaya (myininaya):

\[e^x \ln(e)=\ln(10)\] now ln(e)=1 so we have \[e^x=\ln(10)\]

myininaya (myininaya):

now do ln( ) of both sides again

OpenStudy (anonymous):

is the final answer ln(ln(10)) app. .8340

myininaya (myininaya):

yep! :)